I wrote the problem i wrote before wrong. This is what it really is.

A total of $25,000 is inested in two funds paying 5% and 6% annual interest. The combined annual interest is $1400. How much of te $25,000 is invested in each fund?

Answer by ankor@dixie-net.com(22740)

A total of $25,000 is inested in two funds paying 5% and 6% annual interest. The combined annual interest is $1400. How much of the $25,000 is invested in each fund?
:
let x = amt invested at 6%, then (25000-x) = amt invested at 5%
:
interest at 6% + interest at 5% = 1400
.06x + .05(25000 - x) = 1400
.06x + 1250 - .05x = 1400
.06x - .05x = 1400 - 1250
.01x = 150
x = 150/.01
x = $15,000 invested at 6%
:
25000 - 15000 = $10,000 invested at 5%
:
Check: .05(10000) + .06(15000) = 1400
.